OpenStudy (anonymous):

which of the following is a solution of x^2 + 14x + 112 = 0 (round to the nearest hundredth

OpenStudy (***[isuru]***):

hi, use the formula if ax^2 + bx + c = 0 then \[x =\frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] so in ur question a = 1 b = 14 c = 112 just plug those values in to the formula ... and u could get the solutions easily did u get it ?

OpenStudy (anonymous):

Nooooo :/

OpenStudy (***[isuru]***):

don't u know about the formula that can be use to find the roots of a quadratic expression ?

OpenStudy (anonymous):

nope , i suck at this lol

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (jazzyfa30):

 try answering this